How do u solve this --> 3x+6=2(x-1)
Answer by Here2Help(19)

Here you go:
3x+6=2(x-1)
3x+6=2x-2 -- use distributive property here on the second part of problem
3x-2x+6=2x-2x-2
1x+6=-2
1x+6-6=-2-6
1x=-8
1x/1=-8/1
Final answer: x=-8